Telescope Bootstrap Startup: 15 September 2012 UT

Observer: DMiller (LBTO), OKuhn, MEdwards(Tucson)
Telescope Operator: GBechetti (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: JUrban (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

Details

MODS AGW startup

21:00 Geno switch to build 2012F. Test with an Acquire preset

00:00 Doug tested IRTC image acquisition because IRC/IRS version changed. Had to stop and restart IRTC GUI and then no problems.

00:30 Elliott and Jeff remove retroreflector and fill dewars

01:45 Geno fully opened side vent doors and started mirror ventilation

01:48 Geno retracted M3 on left side. Authorized MODS1/NONE.

02:00 Tip over and start to open chamber doors.

02:12 Wind bumping up to 18 mph. Closed side vent doors.

02:14 Doors open up to Zenith.

02:30 preset in acquire M5_0738

03:10 preset to M5_1568 No other bright stars around it. ptspiral for 20+ steps, but no star

03:15 ptreset

03:20 Olga and Michelle take Mods image. Bright star down and to the right

03:25 IE 60 (from -30), MODS image, move down and right

03:28 IE 30, CA -30 (from 0) move back up and left, and then down and left for CA. Not far enough, mis-calculated

03:32 IE -120, CA 90 to far up and left

03:35 IE -90, CA 60 preset in acquire on-axis. M5 star in center

03:46 preset BS9183 acquire on-axis. On image. down and right.
  • set secondary global z to -0.7 (from -1.0) for a better focus
  • Added +3000 coma on secondary to improve coma (maybe)

03:55 centered up BS9183 on-axis IE = -86, Ca=84

03:58 preset active BS9183 GS=0, 7 iterations and collimated. Guider ~1.5" fwhm. Not bad for first night open and primary 7C out of equilibrium. On-axis guiding and act opt work well. Pupil centered

04:05 preset active BS9183 GS=9, no collimation changes. well collimated. Off-axis guiding and act opt work well
  • IE = -86, Ca=84
  • Secondary, 2.971, -1.56, 1.216, 496, 612, 0
  • Primary, 1.19, -0.46, 0.27, 12.4, 37.5

04:10 Turn telescope over to Olga for MODS testing

MODS startup

04:15 sieveSnap r=6, b=2

04:20 send active preset to StoneO_1136 _/home/lbto/kuhn/MODS/StartupScripts/AlignCheck/StoneO_1136_act0.img 2.3" seeing

04:25 start series of track presets at PA=0,180,90,270. StoneO_1136_trk0.img (b=8,r=3), trk180.img (b=9,r=4), trk90.img (b=10,r=5), trk270 (b=11,r=6). By-eye looks good --- star at center stays at center. I'll try to measure this more carefully - adapting the LBTtools.LBC.rotcenter task (and/or drawing lines...) - later.

04:43 5-dice xy on StoneO_1136 --- we're collimated but seeing is awful...1.9-2". Airmass 1.205
  • r: 7-11
  • b: 12-16

05:00 5-dice (ra,dec) on StoneO_1136 ... seeing has improved to <~1.5"... well, up to 1.7" now. RMS WFE is not going much below 450nm - hovering around 430-480 nm. Going to start observation.
  • r: 12-16
  • b: 17-21

05:13 acqMODS feige110.acq. Geno says it is clear overhead --- at sunset some clouds in the valley but there were none overhead and when he looked outside just a while ago, it looked clear and he could see the MW.
  • acqMODS feige110.acq
  • modsAlign -r mods1r.20120915.0017.fits send offset of -0.442 5.519
    • took confirmatory image, 18 (pretty well centered - but seeing variations are causing some scatter)
  • execMODS feige110.obs shell RIPed after first exposure of this script. Paused (by typing pause), however the shell restart caused the preset to be canceled. Resumed - it took another exposure in r - or just readout the first one again? need to check - 120906M8.06p.fits
    • Shell RIP probably caused by high wind speeds.
    • created a recovery script. Used it, but found star a little off (r 21)...our initial acquisition put it a little off too (attributing that to seeing variations). dx = 0.5" given. Confirmatory image (r 22).
    • spectra start at b=23, r=23. Seeing is not very good still ~1.7-2". Airmass is 1.333.

06:03 paused script while we examine WFSC images with Doug --- diimples in upper right - is it M1, M2 or readout issues? See, e.g. /Repository/20120915/GCS/left_wfscimage000080.fits and left_wfscimage000111.fits. wfscimage000053.fits.

* realized that we cannot pause a script ... can pause an exposure, but not a script between exposures. I stopped the script.

06:15 Doug reflatted mirror - secondary looked fine. We see the same grey patches as before --- is this a problem to do with the camera or just mirror seeing?

06:21 made a new script for fe110 to use a different guide star (21:19:54.658 -05:14:46.23). fe110_newgs.acq (takes r 29). On the wfs we see the same sort of pattern - dimples in the upper right, see, e.g. left_wfscimage000161.fits.

06:27 Command red go 1 timed out after 120 seconds - retry said red GO exposure already in progress, however exposure countdown bar is at start, 2s... "red go 1" hung on erasing ... I refreshed the GUI and then we "aborted" the exposure, though it was not in the middle of one, and this worked.

06:34 Geno & Doug are clearing forces on M1. This new gs was close to bottom edge of guider field and we see 1 more column of vignetting on the left of the wfs image.

06:42 acqMODS feige110.acq. Guide star is way out of focus. I edited the script to put the guide star on axis - copied target coords to guide coords. Sent preset. We got an error but we see:
Sat Sep 15 06:57:17.656 2012 Preset command failed for left MODS mode ACTIVE with guidestars, Sound=presetFailed.wav
Sat Sep 15 06:57:18.187 2012 Unauthorized Preset command received from both LBTI at bentGregorianCenter focus
Collimated

07:11 recollimated and continuing the observation on Feige110 (prism modes). Airmass 1.267. Still clear. Seeing 1.8". The wfs images still show a similar pattern of "dimples" in the upper right and elongated images at the bottom. See, left_wfscimage000221.fits for an example.

07:30 Blue hung at Erasing the CCD... script didn't time out like it did earlier tonight when this happened. Refreshed GUI, clicked abort, then refreshed again. Seemed to be still hung, so I cntl"c'ed script:
ERROR: Command 'abort' timed out after 30 seconds ... abort sequence terminated with errors.
Did "blue reset" and then ran script "execMODS -b Blueprism feige110prism.obs". Took images 29,30,31.

07:48 MODS checkout completed. switching to Michelle for LBC testing

07:55 Noticed that winds on secondary were very high (10-11 m/s) while at zenith and switching instruments. Winds out of the north east are 17 m/s and acceptable but obviously we'll stay out to the west to avoid these high winds on the secondary.

08:14 Slewing to first LBC binocular preset, WT10_364, for pointing, focos, collimation check

08:33 Red side not converging. Trying an initialize. Redoing dofpia

08:45 Adding 1000 nm of z11; no appreciable change. Noting that temperatures of the mirror on the right side is 6.85C and ambient is 3.62C and the left side is 4.19C and ambient is 3.35 C. This could be a temperature issue

08:49 Zeroing out DX mirror again and adding z11 back. Taking image to see if this helps.

08:55 Taking extra and intra focal OB to see if FPIA is putting us on the wrong side of focus. Note that we're at summer position and the ambient temp is 3C; blue is at the lower limit of z for extrafocal and red is at -1.4.

09:10 Adding another 1000 in z11 and -100 in z22

09:20 dofpia converged

09:30 lbcrangebal - ran correctly

09:45 Running rotation center check with rotatorcenter_pa0 and rotatorcenter_pa180 OBs

09:49 Presetting to new, dense field for guiding check

09:58 dofpia

10:00 While doing analysis of rotator center, we realize that the PA for the PA = 180 image was actually slightly out of range for the rotcenter iraf task; redoing that check in new field position

10:20 Taking longer exposures to understand rotator position as written to header and exercising instrument

10:40 Continued longer exposure for guiding. Confirm that Tech Chip focus corrections are being calculated and sent. Checked that focus corrections are being received by PCS

10:55 Slew to USNO star USNOB1300-0063601 for rangebal, field rotation, and dither check

10:58 dofpia -- ran into focus limits on SX side; continuing

11:06 rangebal --

11:10 Running test of dither, field rotation. Dither pattern: 10, 15; 25, 30; 45, 45; 45, 20

11:32 Preset to dense field for Focal Plane Image Quality

11:35 dofpia

11:43 run RB_rVexin4, RB_rVinex4_pa180,

11:50 dofpia

11:55 Superfoc sequence

12:12 Preset to blank sky for twilight flats

12:35 close - slewing to horizon

-- DougMiller - 14 Sep 2012
Topic revision: r4 - 15 Sep 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback